# Convergence Analysis of Non-Strongly-Monotone Stochastic Quasi-Variational Inequalities

### v3

@@ 6 Numerical Experiment
Over-parameterized Regression Game. In a regression problem, the goal is to find a parameter vector $x\in\mathbb{R}^{d}$ that minimizes the loss function $\ell^{\text{tr}}(x)$ over the training dataset $D^{\text{tr}}$. Without explicit regularization, an over-parameterized regression problem exhibits multiple global minima over the training dataset, and not all optimal regression coefficients perform equally well. Considering a secondary objective, such as minimizing the loss over a validation set $D^{\text{val}}$, helps in selecting a model parameter that performs well on both training and validation dataset.
Figure 1: Comparison of iG-SQVI and iEG-SQVI: (Top) for the triazines dataset with 180 data points and 60 features, (Middle) for the eunite2001 dataset with 320 data points and 16 features, (Bottom) for a synthetic dataset with 250 data points and 25 features.
Consider a collection of $N$ players each having a model parameter $x_{i}\in\mathbb{R}^{d}$. Define $\mathbf{x}\triangleq[x_{i}]_{i=1}^{N}$, and suppose there is a shared training dataset $D^{\text{tr}}$ and each player possesses an individual validation dataset $D^{\text{val}}_{i}$. The goal is to find a model parameter $\mathbf{x}$ by minimizing the training loss $\ell^{\text{tr}}(x)$ while each player improves its model parameter based on their validation set by minimizing $\ell^{\text{val}}_{i}(x_{i})$. This problem can be formulated as the following bilevel GNE:
$\min_{x_{i}\in\mathbb{R}^{d}}\ \ell^{\text{val}}_{i}(x_{i}),\ \mbox{such that}\ x_{i}\in\mbox{argmin}_{x_{i}\in K_{i}}\ell^{\text{tr}}\left(x_{i},x_{(-i)}^{*}\right).$
In this experiment, we define $\ell^{\text{val}}_{i}(x_{i})\triangleq\frac{1}{2}\|A_{i}^{\text{val}}x_{i}-b_{i}^{\text{val}}\|^{2},$ where $A_{i}^{\text{val}}\in\mathbb{R}^{n\times d}$ and $b_{i}^{\text{val}}\in\mathbb{R}^{n\times 1}$, and $\ell^{\text{tr}}(x)\triangleq\frac{1}{2}\|A^{\text{tr}}\mathbf{x}-b^{\text{tr}}\|^{2},$ where $A^{\text{tr}}\in\mathbb{R}^{Nn\times Nd}$ and $b^{\text{tr}}\in\mathbb{R}^{Nn\times 1}$ and $X_{i}=\{x_{i}\mid\|x_{i}\|\leq\lambda\}$ for some $\lambda>0$. One can show that this problem can be formulated as (SQVI) by choosing $K(x)=\Pi_{i=1}^{N}K_{i}(x_{(-i)})$, where
$K_{i}(x_{(-i)})=\mbox{argmin}_{x_{i}\in X_{i}}\frac{1}{2}\left\|A_{i}^{\text{tr}}x_{i}+A_{(-i)}^{\text{tr}}x_{(-i)}-b_{i}^{\text{tr}}\right\|^{2},$
@@
$F_{i}(x_{i})=(A_{i}^{\text{val}})^{T}(A_{i}^{\text{val}}x_{i}-b_{i}^{\text{val}}).$
Note that, operator $F$ is monotone and satisfies quadratic growth property but it may not be strongly monotone. Furthermore, one can show that Assumption 3 is satisfied (see Theorem 8 in [32] and Corollary 2 in [5]). In Figure 1, we present a performance comparison of our proposed methods. For the triazines dataset [9], we set the number of players $N=6$, for the eunite2001 dataset [9] $N=4$, and for the synthetic dataset $N=10$. In all cases, we utilized $80\%$ of the data points for training and allocated the remaining $20\%$ for validation. To solve the projection inexactly, observe that the sub-problem is a simple bilevel optimization problem. This type of problem has been explored in the literature [21, 44]. Here, following [44], we employed the FISTA algorithm [4] to solve the corresponding regularized problem satisfying Assumption 4. For all the experiments, we execute the inner algorithm for $k\log^{2}(k+1)(1-1\text{e-}3)^{k}$ iterations, and the remaining parameters are selected according to the following table after fine-tuning.
@@
triazines eunite2001 synthetic
Stepsize $\eta$ 5e-2 3e-1 1e-2
$\bar{\alpha}$ 1e-1 5e-1 9e-1
$\bar{b}$ 1e-1 5e-1 12e-1
Regularizer 1e0 1e-1 1e-2
@@ 7 Conclusion
Acknowledgments
This work is supported in part by the National Science Foundation under Grant ECCS-2231863, the University of Arizona Research, Innovation & Impact (RII) Funding, and the Arizona Technology and Research Initiative Fund (TRIF) for Innovative Technologies for the Fourth Industrial Revolution initiatives.
@@ Appendix A Extra Gradient Method
$\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\bar{\alpha}(1-q)^{T-1}\sum_{k=0}^{T-1}\left(\frac{\eta\nu}{(1-q)^{k}\sqrt{N}}+\mathbb{E}[\|e_{k}\|](1-q)^{-k}\right).}$
@@
where in the last inequality we used the definition of $T$ and $N$$\square$
@@
$\quad+{\bar{\alpha}(1-q)^{T-1}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{E}[\|e_{k}\|]}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1-q)}^{-k}\right).}$
@@ Appendix B Gradient Method
$\quad+\alpha_{k}\|\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]$
$\qquad-\mathcal{P}_{K(\bar{x}_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]\|$
$\quad+\alpha_{k}\|\mathcal{P}_{K(\bar{x}_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]$
@@
$\quad+\alpha_{k}\underbrace{\|x_{k}-\bar{x}_{k}-\eta(F(x_{k})-F(\bar{x}_{k}))\|}_{\text{term (a)}}+\alpha_{k}\eta\|\bar{w}_{k,N_{k}}\|+{\alpha_{k}\|e_{k}\|.}$ (A2)
@@
$\quad-2\eta\langle x_{k}-\bar{x}_{k},F(x_{k})-F(\bar{x}_{k})\rangle$
$\leq(1+L^{2}\eta^{2}-2\eta k_{f})\|x_{k}-\bar{x}_{k}\|^{2}$
[…]
Now by using (B) in (B), defining $\beta\triangleq\gamma+\sqrt{1+L^{2}\eta^{2}-2\eta\mu}$ and $q_{i}\triangleq(1-\beta)\alpha_{i}$ we get the following:
@@
$\quad(1-\alpha_{k})\|x_{k}-\bar{x}_{k}\|+\alpha_{k}\eta\|\bar{w}_{k,N_{k}}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\alpha_{k}\|e_{k}\|}$
$\qquad+\alpha_{k}\left(\gamma+\sqrt{1+L^{2}\eta^{2}-2\eta\mu}\right)\|x_{k}-\bar{x}_{k}\|$
@@
$\qquad+\sum_{i=0}^{k-1}\left(\left(\prod_{j=i}^{k-1}(1-q_{j+1})\right)\alpha_{i}\left(\eta\|\bar{w}_{i,N_{i}}\|+\|e_{i}\|\right)\right).$
Next, by choosing $\alpha_{k}=\bar{\alpha}$, where $10<\bar{\alpha}<$ 1, and based on the conditions of the theorem, one can easily verify that $\beta<1$ and $q_{k}=q<1$ for all $k\geq 0$. $\square$
@@
$\quad+{\bar{\alpha}(1-q)^{T-1}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{E}[\|e_{k}\|]}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1-q)}^{-k}\right).}$
Since $\rho\geq 1-q$ and $q\in(0,1)$, one can easily show that $-\frac{(1-q)^{T}}{\rho+q-1}<0$. Hence, from the previous inequality, we conclude that
@@
$\quad+{\bar{\alpha}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}\left(\mathbb{E}[\|e_{k}\|]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1-q)^{T-1-k}}\right)}$
$\leq\rho^{T}\|x_{0}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}_{0}}\|+\frac{\bar{\alpha}\eta\nu\rho^{T}}{\rho+q-1}+\bar{\alpha}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}\left(\mathbb{E}[\|e_{k}\|]\rho^{T-1-k}\right).$
@@ Appendix C Almost Sure Convergence
$\quad+\alpha_{k}(e_{k}+\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]-\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right])\|^{2}.$
@@
$\quad+2\langle x_{k}-\bar{x}_{k},\alpha_{k}e^{\prime}_{k}\rangle-2\langle H(x_{k}),\alpha_{k}e^{\prime}_{k}\rangle.$ (A5)
@@
$\langle\eta F(\bar{x}_{k}),\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right]-\bar{x}_{k}\rangle\geq 0.$
@@
$\langle x_{k}-\eta F(x_{k})-\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right],\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right]-\bar{x}_{k}\rangle\geq 0.$
Combining this inequality with the previous one, and using definition of $H(x_{k})$ we get:
$\langle\tfrac{1}{\alpha_{k}}H(x_{k})-\eta(F(x_{k})-F(\bar{x}_{k})),-\tfrac{1}{\alpha_{k}}H(x_{k})+(x_{k}-\bar{x}_{k})\rangle\geq 0.$

### v4

@@ 6 Numerical Experiment
Over-parameterized Regression Game. In a regression problem, the goal is to find a parameter vector $x\in\mathbb{R}^{d}$ that minimizes the loss function $\ell^{\text{tr}}(x)$ over the training dataset $D^{\text{tr}}$. Without explicit regularization, an over-parameterized regression problem exhibits multiple global minima over the training dataset, and not all optimal regression coefficients perform equally well. Introducing a secondary objective provides a principled way to select among these solutions: it guides us toward parameters that not only fit the training data but also optimize an additional desirable property. For instance, using the $\ell_{1}$-norm as a secondary objective favors sparse solutions, while using the $\ell_{2}$-norm selects solutions with smaller overall magnitude.
Figure 1: Comparison of iG-SQVI and iEG-SQVI: (Top) for the wikivital mathematics dataset with 300 data points and 1065 features, (Bottom) for a synthetic dataset with 60 data points and 500 features.
Consider a collection of $N$ players each having a model parameter $x_{i}\in\mathbb{R}^{d_{i}}$. Define $\mathbf{x}\triangleq[x_{i}]_{i=1}^{N}$, and suppose there is a shared training dataset $D^{\text{tr}}$ and the goal is to find a model parameter $\mathbf{x}$ by minimizing the training loss $\ell^{\text{tr}}(\mathbf{x})$ while each player improves its model parameter based on their secondary objective by minimizing a weighted $\ell_{2}$-norm. This problem can be formulated as the following bilevel GNE:
$\min_{x_{i}\in X_{i}}\ \ell^{W}_{i}(x_{i})\quad\mbox{s.t.}\quad x_{i}\in\mbox{argmin}_{x_{i}\in\mathbb{R}^{d}}\ell^{\text{tr}}\left(x_{i},x_{(-i)}^{*}\right).$
In this experiment, we define $\ell^{\text{tr}}(\mathbf{x})\triangleq\frac{1}{2}\|A^{\text{tr}}\mathbf{x}-b^{\text{tr}}\|^{2},$ where $A^{\text{tr}}\in\mathbb{R}^{n\times d}$, $d=\sum_{i=1}^{N}d_{i}$, $n<\min_{i}\{d_{i}\}$, $b^{\text{tr}}\in\mathbb{R}^{n\times 1}$, and $\ell^{W}_{i}(x_{i})\triangleq\frac{1}{2}\|W_{i}x_{i}\|^{2},$ where $W_{i}\in\mathbb{R}^{p_{i}\times d_{i}}$ is generated randomly such that $p_{i}<d_{i}$, and $X_{i}=\{x_{i}\mid\|x_{i}\|_{1}\leq\lambda\}$ for some $\lambda>0$. One can show that this problem can be formulated as QVI by choosing $K(x)=\Pi_{i=1}^{N}K_{i}(x_{(-i)})$, where
$K_{i}(x_{(-i)})=\mbox{argmin}_{x_{i}\in\mathbb{R}^{d_{i}}}\frac{1}{2}\left\|A_{i}^{\text{tr}}x_{i}+A_{(-i)}^{\text{tr}}x_{(-i)}-b_{i}^{\text{tr}}\right\|^{2},$
@@
$F_{i}(x_{i})=W_{i}^{T}(W_{i}x_{i}).$
Note that, this problem belongs to the class of problem we defined in subsection 4.2. Operator $F$ is monotone and satisfies quadratic growth property but it is not be strongly monotone. Furthermore, one can show that Assumption 3 is satisfied (see Corollary 2 in [5]). In Figure 1, we present a performance comparison of our proposed methods. For the wikivital mathematics dataset [41], we set the number of players $N=3$ and for the synthetic dataset $N=5$. To solve the projection inexactly, observe that the sub-problem is a simple bilevel optimization problem. This type of problem has been explored in the literature [20, 44]. Here, following [44], we employed the FISTA algorithm [4] to solve the corresponding regularized problem satisfying Assumption 4. For all the experiments, we execute the inner algorithm for $k\log^{2}(k+1)(1-10^{-4})^{-k}$ iterations, and the remaining parameters are selected according to the following table after fine-tuning.
@@
wikivital mathematics synthetic
Stepsize $\eta$ 1e-4 1e-4
$\bar{\alpha}$ 5e-2 1e-1
$\bar{b}$ 2.5e-2 1e-1
Regularizer 1e-4 1e-4
@@ 7 Conclusion
Author contribution The manuscript was mainly written and edited by ZA. AJ provided research conditions and guidance. The technical implementations, such as coding algorithms and setting up reproducible environments were done by ZA. All authors read and approved the final manuscript.
Funding This work is supported in part by the National Science Foundation under Grant ECCS-2231863, the University of Arizona Research, Innovation & Impact (RII) Funding, and the Arizona Technology and Research Initiative Fund (TRIF) for Innovative Technologies for the Fourth Industrial Revolution initiatives.
Availability of data and materials The basic code of this work are available from the corresponding author upon reasonable request.
@@ Appendix A Extra Gradient Method
$\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\bar{\alpha}(1-q)^{T-1}\sum_{k=0}^{T-1}\left(\frac{\eta\nu}{(1-q)^{k}\sqrt{N}}+\mathbb{E}[\|e_{k}\|](1-q)^{-k}\right).}$
@@
[…]
@@
$\qquad+{\bar{\alpha}(1-q)^{T-1}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{E}[\|e_{k}\|]}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1-q)}^{-k}\right).}$
@@ Appendix B Gradient Method
$\qquad+{\alpha_{k}e_{k}}-(1-\alpha_{k})\bar{x}_{k}-\alpha_{k}\mathcal{P}_{K(\bar{x}_{k})}\left[\bar{x}_{k}-\eta F(\bar{x}_{k})\right]\|$
@@
$\qquad+\alpha_{k}\|\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]$
$\hskip 18.49988pt-\mathcal{P}_{K(\bar{x}_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]\|$
$\qquad+\alpha_{k}\|\mathcal{P}_{K(\bar{x}_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]$
$\hskip 18.49988pt-\mathcal{P}_{K(\bar{x}_{k})}\left[\bar{x}_{k}-\eta F(\bar{x}_{k})\right]\|+{\alpha_{k}\|e_{k}\|}$
@@
$\qquad+\alpha_{k}\underbrace{\|x_{k}-\bar{x}_{k}-\eta(F(x_{k})-F(\bar{x}_{k}))\|}_{\text{term (a)}}+\alpha_{k}\eta\|\bar{w}_{k,N_{k}}\|+{\alpha_{k}\|e_{k}\|.}$ (A2)
@@
$\qquad-2\eta\langle x_{k}-\bar{x}_{k},F(x_{k})-F(\bar{x}_{k})\rangle$
$\leq(1+L^{2}\eta^{2}-2\eta\mu_{F})\|x_{k}-\bar{x}_{k}\|^{2}$
$\implies\text{term(a)}\leq\sqrt{1+L^{2}\eta^{2}-2\eta\mu_{F}}\|x_{k}-\bar{x}_{k}\|.$ (A3)
Now by using (B) in (B), defining $\beta\triangleq\gamma+\sqrt{1+L^{2}\eta^{2}-2\eta\mu_{F}}$ and $q_{i}\triangleq(1-\beta)\alpha_{i}$ we get the following:
@@
$\qquad(1-\alpha_{k})\|x_{k}-\bar{x}_{k}\|+\alpha_{k}\eta\|\bar{w}_{k,N_{k}}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\alpha_{k}\|e_{k}\|}$
$\hskip 18.49988pt+\alpha_{k}\left(\gamma+\sqrt{1+L^{2}\eta^{2}-2\eta\mu_{F}}\right)\|x_{k}-\bar{x}_{k}\|$
@@
$\hskip 18.49988pt+\sum_{i=0}^{k-1}\left(\left(\prod_{j=i}^{k-1}(1-q_{j+1})\right)\alpha_{i}\left(\eta\|\bar{w}_{i,N_{i}}\|+\|e_{i}\|\right)\right).$
Next, by choosing $\alpha_{k}=\bar{\alpha}$, where $0<\bar{\alpha}<1$, and based on the conditions of the theorem, one can easily verify that $\beta<1$ and $q_{k}=q<1$ for all $k\geq 0$. $\square$
@@
$\qquad+{\bar{\alpha}(1-q)^{T-1}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{E}[\|e_{k}\|]}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1-q)}^{-k}\right).}$
Since $\rho\geq 1-q$ and $q\in(0,1)$, one can easily show that $\frac{(1-q)^{T}}{\rho+q-1}<0$. Hence, from the previous inequality, we conclude that
@@
$\qquad+{\bar{\alpha}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}\left(\mathbb{E}[\|e_{k}\|]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1-q)^{T-1-k}}\right)}$
$x$ $\leq\rho^{T}\|x_{0}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}_{0}}\|+\frac{\bar{\alpha}\eta\nu\rho^{T}}{\rho+q-1}+\bar{\alpha}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}\left(\mathbb{E}[\|e_{k}\|]\rho^{T-1-k}\right).$
@@ Appendix C Almost Sure Convergence
$\qquad+\alpha_{k}(e_{k}+\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]-\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right])\|^{2}.$
@@
$\qquad+2\langle x_{k}-\bar{x}_{k},\alpha_{k}e^{\prime}_{k}\rangle-2\langle H(x_{k}),\alpha_{k}e^{\prime}_{k}\rangle.$ (A5)
@@
$\langle\eta F(\bar{x}_{k}),\mathcal{P}_{K(\bar{x}_{k})}\left[x_{k}-\eta F(x_{k})\right]-\bar{x}_{k}\rangle\geq 0.$
@@
$\langle x_{k}-\eta F(x_{k})-\mathcal{P}_{K(\bar{x}_{k})}\left[x_{k}-\eta F(x_{k})\right],\mathcal{P}_{K(\bar{x}_{k})}\left[x_{k}-\eta F(x_{k})\right]-\bar{x}_{k}\rangle\geq 0.$
Defining $\widehat{H}(x_{k})\triangleq\alpha_{k}(x_{k}-\mathcal{P}_{K(\bar{x}_{k})}\left[x_{k}-\eta F(x_{k})\right])$ and combining this inequality with the previous one, we get:
$\langle\tfrac{1}{\alpha_{k}}\widehat{H}(x_{k})-\eta(F(x_{k})-F(\bar{x}_{k})),-\tfrac{1}{\alpha_{k}}\widehat{H}(x_{k})+(x_{k}-\bar{x}_{k})\rangle\geq 0.$